OpenStudy (anonymous):

????? Which statement is true about the data? 10, 12, 9, 12, 8, 10, 11, 8 The range is 1. The median is 10. The mean is 9. There are 2 modes.

OpenStudy (anonymous):

I will give medal and be fan

OpenStudy (anonymous):

to get the average we need to add all the values together and divide that by the number of numbers: 10 + 12 + 9 + 12 + 8 + 10 + 11 + 8 = 80 and there are 8 numbers so the mean (or the average) = 80/8 = 10 which does not concur with the possible answer The median will be the number that numerically separates the lower half of the values from the upper half. arranging the numbers in order of magnitude 8 8 9 10 10 11 12 12 As there are eight numbers and eight is an even number, there will be two median numbers. from the above it is clear that they are ten and ten. to get a single median value we add them them and divide by the number of median values: 10 +10 = 20 and 20/2 = 10 So ten is the indeed the median value Note that there are more than two modes: 8, 10 and 12 all occur twice and thus are all modes of this group of numbers. To find the range you subtract the smallest number from the largest number: 12 - 8 = 4